OpenStudy (anonymous):

whats the answer for 1 srqt of 1+260pi^2 all over 2 pi

OpenStudy (anonymous):

1 + srqt of 1+260pi^2 all over 2 pi

OpenStudy (turingtest):

you want to simplify\[\LARGE\frac1{\sqrt{1+260\pi^2}}\over2\pi\]correct?

OpenStudy (anonymous):

idk

Directrix (directrix):

Consider clicking on the blue "DRAW" button and manually writing the expression to be simplified. Thanks.

OpenStudy (anonymous):